You determined the molar mass of an unknown fatty acid 3 dissolved into stearic acid to be 219.4 g/mol. Unknown fatty acid 3 was myristic acid, which has a molar mass of 228.37 g/mol. What was your percent error to the correct number of significant figures?
Answer #1

Percent error is calculated as:

Difference of Experimental value (E) and Theoretical value (T) , divided by T , is multiplied by 100.

Here E = 219.4 g/mol , T = 228.37 g/mol

Therefore the Difference in values = T - E = (228.37 - 219.4 )g/mol = 8.97 g/mol

The error = 8.97 / T = 8.97 / 228.37 = 0.0393

The percent error = 0.039278 x 100 % = 3.9278 %

Now, the percent error is always expressed in one or maximum 2 significant digits.

Therefore Rounding the number 3.9278 % to 2 significant digits gives us 0.039 or 3.9 %